\begin{thebibliography}{}
\footnotesize

\bibitem[Aus11]{Austin}Austin, T. (2011). Deducing the density Hales–Jewett theorem from an infinitary removal lemma. Journal of Theoretical Probability, 24(3), 615.

\bibitem[Ber96]{Bergelson2}Bergelson, V.,  Leibman, A. (1996). Polynomial extensions of van der Waerden’s and Szemerédi’s theorems. Journal of the American Mathematical Society, 9(3), 725-753.

\bibitem[Ber99]{Bergelson1}Bergelson, V.,  Leibman, A. (1999). Set-polynomials and polynomial extension of the Hales-Jewett theorem. Annals of Mathematics, 33-75.

\bibitem[Blog09]{Gowersblog} Gowers, W.T.(2009, November 14) The first unknown case of polynomial DHJ
[Blog Post]. Retrieved from https://gowers.wordpress.com/2009/11/14/the-first-unknown-case-of-polynomial-dhj/

\bibitem[Blog21]{Kalaiblog} Kalai, G.(2021, January 29) Possible future Polymath projects (2009, 2021)
[Blog Post]. Retrieved from https://gilkalai.wordpress.com/2021/01/29/possible-future-polymath-projects-2009-2021/

\bibitem[Dod14]{Dodos}Dodos, P., Kanellopoulos, V.,  Tyros, K. (2014). A simple proof of the density Hales–Jewett theorem. International Mathematics Research Notices, 2014(12), 3340-3352.

\bibitem[Fur77]{Furstenberg2} Furstenberg, H. (1977). Ergodic behavior of diagonal measures and a theorem of Szemerédi on arithmetic progressions. Journal d’Analyse Mathématique, 31(1), 204-256.

\bibitem[Fur91]{Furstenberg1} Furstenberg, H., Katznelson, Y. (1991). A density version of the Hales-Jewett theorem. Journal d’Analyse Mathématique, 57(1), 64-119.

\bibitem[Gow01]{Gowers2}Gowers, W. T. (2001). A new proof of Szemerédi's theorem. Geometric and Functional Analysis GAFA, 11(3), 465-588.

\bibitem[Gow04]{Green} Green, B. (2004). Finite field models in additive combinatorics. arXiv preprint math/0409420.

\bibitem[Gow06]{Gowers3}Gowers, W. T. (2006). Quasirandomness, counting and regularity for 3-uniform hypergraphs. Comb. Probab. Comput., 15(1-2), 143-184.

\bibitem[Gow07]{Gowers4}Gowers, W. T. (2007). Hypergraph regularity and the multidimensional Szemerédi theorem. Annals of Mathematics, 897-946.

\bibitem[McC11]{McCutcheon} McCutcheon, R. (2011). The wildcard set's size in the density Hales-Jewett theorem. The Electronic Journal of Combinatorics, P114-P114.

\bibitem[Nag06]{Nagle}Nagle, B., Rödl, V.,  Schacht, M. (2006). The counting lemma for regular k‐uniform hypergraphs. Random Structures and Algorithms, 28(2), 113-179.

\bibitem[Pol12]{Polymath}Polymath, D. H. J. (2012). A new proof of the density Hales-Jewett theorem. Annals of Mathematics, 1283-1327.

\bibitem[Rod04]{Rodl}Rödl, V.,  Skokan, J. (2004). Regularity lemma for k‐uniform hypergraphs. Random Structures and Algorithms, 25(1), 1-42.

\bibitem[Sze75]{Szemeredi}Szemerédi, E. (1975). On sets of integers containing no k elements in arithmetic progression. Acta Arith, 27(199-245), 2.

\bibitem[Hal09]{Hales} Hales, A. W., Jewett, R. I. (2009). Regularity and positional games. In Classic Papers in Combinatorics (pp. 320-327). Birkhäuser Boston.

\bibitem[Sar78]{Sarkozy}Sárközy, A. (1978). On difference sets of sequences of integers. I. Acta Mathematica Academiae Scientiarum Hungarica, 31(1-2), 125-149.

\bibitem[van27]{Waerden}van der Waerden, B. (1927). Beweis einer baudetschen vermutung. Nieuw Arch. Wisk., 19, 212-216.

\bibitem[Wal00]{Walters} Walters, M. (2000). Combinatorial proofs of the polynomial van der Waerden theorem and the polynomial Hales–Jewett theorem. Journal of the London Mathematical Society, 61(1), 1-12.








\end{thebibliography}

\documentclass[letter,11pt]{article}

\newcounter{conj}
\usepackage{pifont}

\usepackage{comment}
\usepackage[linktocpage=true]{hyperref}
\usepackage[margin=1.2 in, top=1 in, bottom= 1.2 in]{geometry}

\usepackage[english]{babel}
\usepackage{amsfonts}
\usepackage{mathrsfs}
\usepackage[mathscr]{euscript}
\usepackage{bbm}
\usepackage{latexsym}
\usepackage{math dots}
\usepackage{amssymb}
\usepackage{mathtools}

\usepackage{enumitem}
\setlist{nolistsep}
\usepackage{amsthm}
\usepackage{relsize}
\usepackage{nicefrac}
\usepackage[capitalize]{cleveref}

\newtheoremstyle{plain}{3mm}{3mm}{\slshape}{}{\bfseries}{.}{.5em}{}
\newtheoremstyle{definition}{2mm}{2mm}{}{}{\bfseries}{.}{.5em}{}
\theoremstyle{plain}
	
\newtheorem{theorem}{Theorem}

\newtheorem{conjecture}[conj]{Conjecture}

\theoremstyle{definition}

\theoremstyle{plain}
\newtheorem*{namedthm}{\namedthmname}
\newcounter{namedthm}
	

\newcommand{\N}{\mathbb{N}}
\newcommand{\Z}{\mathbb{Z}}

\title{Difference sets in quadratic density Hales Jewett conjecture with 2 letters}
\author{Aritro Pathak}

\begin{document}

\maketitle

\begin{abstract}
 The Quadratic Density Hales Jewett conjecture with $2$ letters states that for large enough $n$, every dense subset of $\{0,1\}^{n^{2}}$ contains a combinatorial line where the wildcard set is of the form $\gamma \times \gamma$ where $\gamma \subset \{1,2,\dots n\}$. We show in an elementary quantitative way that every dense subset of $\{0,1\}^{n^{2}}$, for sufficiently large $n$,  contains two elements such that the set of coordinate points where they differ, which we term the difference set of these two elements, is of the form $\gamma_{1}\times \gamma_{2}$ where $\gamma_1, \gamma_2$ are both nonempty subsets of $\{1,2,\dots n\}$. Further we give several non-trivial examples of dense vector subspaces of $\{0,1\}^{n^{2}}$, where in each case the wildcard set of the combinatorial line that can be obtained has restrictions on its size and shape.
\end{abstract}

\section{Introduction}

Consider for any $n\in \Z_+$, the $n \times n$ grid where each entry $\{(i,j):1\leq i,j\leq n \}$ is filled with either 0 or 1, i.e the set $\{0,1\}^{n^{2}}$.

The following is the first case of a central open conjecture in Ramsey theory, as mentioned by Gowers\cite{Gowersblog}, and Bergelson. This question has been brought up again with renewed interest in a recent blog post by Kalai\cite{Kalaiblog}.

\begin{conjecture}[Quadratic Density Hales Jewett with 2 letters] \label{conj:conj1}
For any $0<\delta<1$, there exists $QDHJ(\delta)$ so that for any $n\geq QDHJ(\delta)$, for any subset $S\subset \{0,1\}^{n^2}$ with $|S|\geq \delta \cdot 2^{n^2}$, there exist two elements $s^{(0)}, s^{(1)}\in S$ such that the set $\{(i,j):s^{(0)}_{ij}\neq s_{ij}^{(1)}\}$ is the same where $\{ s^{(0)}_{ij}=0,s^{(1)}_{ij}=1\}$, where $\{(i,j)\in \gamma\times \gamma, \ \gamma\subset \{1,2,..,n\}\}$ with $\gamma$ nonempty. 
\end{conjecture}

This is the quadratic base case with $k=2$ letters of the general Polynomial Density Hales Jewett conjecture stated later as \cref{conj:PDHJ}. To state this conjecture formally, we first introduce some notation, which we essentially borrow from Walters \cite{Walters}.

For any given $n\in \mathbb{Z}_{+}$, consider the set of words of length $n$, with each letter of the word being an element of $[k]:=\{1,2,..,k\}$. Formally this is written as $K=[k]^{n}$, and we also denote the set $\{1,2,\dots,n\}$ by $[n]$. For any $a\in K, \gamma \subset [n]$ and $1\leq x\leq k$ we denote by $b=a\oplus x\gamma$ the element of $[k]^{n}$ which is obtained by setting $b_i=x$ if $i \in \gamma$ and $b_i=a_i$ otherwise. A \textit{combinatorial line} is a set of the form $\{a\oplus x\gamma :1\leq x\leq k\}$. We call $\gamma$ the wildcard set for the combinatorial line. A \textit{k-dimensional combinatorial subspace} is a set of the form $\{a\oplus x_1 \gamma_1 \oplus x_2 \gamma_2 \dots \oplus x_k \gamma_k : 1\leq x_{i}\leq k, \forall 1\leq i\leq k \}$

In the context of the Polynomial Hales Jewett theorem, we are looking at not just a linear coordinate space such as $\{1,2,\dots,n\}$ but at $d$-dimensional coordinate grids, $\{1,2,\dots,n\}^{d}$ for all positive integers $d$. In these cases, for the set of words of length $n^d$ with $k$ letters, we use the symbol $[k]^{n^{d}}$. When $k=2$, for any two words $s_1,s_2 \in \{0,1\}^{n^{d}}$, we call the \textit{difference set} of $s_1,s_2$ the set of coordinate points where the words differ. 

We state the Polynomial Hales Jewett theorem in a form articulated by Walters \cite{Walters}, and later by McCutcheon\cite{McCutcheon}, and which is a generalization of the original theorem proven by Bergelson and Leibman\cite{Bergelson1}.

\begin{theorem}[Polynomial Hales Jewett:]\label{thm:polyhales}
 For positive integers $k,r$, there exists a positive integer $N(k,r)$ such that for all $n\geq N(k,r)$ whenever $K[n]:=[k]^{n} \times [k]^{n^{2}}\times \dots \times [k]^{n^{d}}$ is $r$ colored, there exists $a\in K[n]$ and $\gamma\subset [N]$ such that the set of points $\{a\oplus x_1\gamma \oplus x_2 (\gamma \times \gamma)\oplus \dots \oplus x_{d} \gamma^{d}:1\leq x_{i} \leq k\}$ is monochromatic.
\end{theorem}

As noted by Walters\cite{Walters}, the original Polynomial Van der Waerden theorem follows straightforwardly from this Polynomial Hales Jewett theorem. 

Stronger density versions of these coloring statements have been established over the years. The density version of van der Waerden's theorem\cite{Waerden} is Szemeredi's theorem, of which several celebrated proofs are known. \cite{Szemeredi,Furstenberg2,Gowers2, Gowers3, Gowers4, Nagle, Rodl}.

\begin{theorem}[Szemeredi:]\label{thm:szemeredi}
For every positive integer $k$ and every $\delta>0$ there exists a positive integer $N(k,\delta)$ so that for any $n\geq N(k,\delta)$ and every subset $A\subset [n]$ of size at least $\delta\cdot n$ contains an arithmetic progression of length $k$.
\end{theorem}

It is easy to show that the original Hales Jewett theorem\cite{Hales} implies the classical van der Waerden theorem by considering the base $k$ representations of the integers in question. In 1991, Furstenberg and Katznelson proved the density version of Hales Jewett theorem \cite{Furstenberg1} from which Szemeredi's theorem follows as a corollary. Later several other proofs were also found\cite{Polymath, Austin, Dodos}.

We state the Density Hales Jewett theorem here:

\begin{theorem}[Density Hales Jewett:]\label{thm:densityhales}
For every positive integer $k$ and any $0<\delta<1$, there exists a positive integer $DHJ(k,\delta)$ such that for any $n\geq DHJ(k,\delta)$ , a subset $A\subset[k]^{n}$ with density at least $\delta$ contains a combinatorial line of size k.
\end{theorem}

The $k=2$ version of the density Hales Jewett theorem follows from Sperner's theorem; where we interpret a string of $0'$s and $1'$s of length $n$ to be a subset of $\{1,2,...,n\}$, and a combinatorial line is interpreted as two subsets $A,B$ such that one of them, say $A$ is completely contained inside the other. Thus to not contain a combinatorial line, we are looking at a maximal antichain which by Sperner's theorem has size at most ${{n}\choose {\lfloor \frac{n}{2} \rfloor}}$, and by Stirling's approximation, $\frac{{{n}\choose {\lfloor \frac{n}{2} \rfloor}}}{2^{n}}\sim \frac{2}{\sqrt{n}} \to 0$ as $n\to \infty$. Thus for large enough $n$, we are forced to contain a combinatorial line.

Later, Bergelson established the polynomial van der Waerden theorem and polynomial Szemeredi theorem in the same paper\cite{Bergelson2}, for polynomials with integer coefficients and without constant terms. Such polynomials are refered to as integral polynomials. We state the theorems below:

\begin{theorem}[Polynomial van der Waerden:]\label{thm:polyvander}
Suppose that $p_1,p_2,..,p_m$ are $m$ integral polynomials. Then there exists a positive integer $N(m)$ such that for all $n\geq N(m)$, there exists $a$ and $0<d\leq n$ such that the set $\{a\} \cup \{a+p_i(d):1\leq i\leq m\}$ is monochromatic, whenever $\N$ is finite colored.

\end{theorem}

\begin{theorem}[Polynomial Szemeredi:]\label{thm:polyszemeredi}

Suppose that $p_1,p_2,..,p_m$ are $m$ integral polynomials and $0<\delta<1$ any small real number. Then there exists a positive integer $N(m, \delta)$ such that for all $n\geq N(m,\delta)$, and any set $A\subset [n]$ of size at least $\delta\cdot n$, there exists $a$ and $0<d\leq n$ such that the set $\{a\} \cup \{a+p_i(d):1\leq i\leq m\}$ is monochromatic.

\end{theorem}

There is a natural polynomial density Hales Jewett conjecture that generalizes \cref{thm:polyszemeredi}, which we state below. It would be a generalization of both the Polynomial density Hales Jewett theorem as well as the Density Hales Jewett theorem. For $k=2$, and only considering the quadratic term, we stated the corresponding version in \cref{conj:conj1}.

\begin{conjecture}[Polynomial Density Hales Jewett]\label{conj:PDHJ}

For positive integers $k$ and a small $0<\delta<1$, there exists a positive integer $N(k,\delta)$ such that for all $n\geq N(k,\delta)$ whenever we consider $A\subset K[n]=[k]^{n} \times [k]^{n^{2}}\times \dots \times [k]^{n^{d}}$ with cardinality at least $\delta\cdot k^{n+n^{2}+\dots+n^{d}}$, there exists $a\in K[n]$ and $\gamma\subset [N]$ such that the set of points $\{a\oplus x_1\gamma \oplus x_2 (\gamma \times \gamma)\oplus \dots \oplus x_{d} \gamma^{d}:1\leq x_{i} \leq k, 1\leq i\leq d \}$ is contained in $A$.
\end{conjecture} 


Here we briefly outline how the conjectured Quadratic Density Hales Jewett theorem implies Sarkozy's theorem\cite{Sarkozy} which is a special case of the Polynomial Szemeredi theorem outlined above. By a similar argument to the one outlined below, one can also show that the Polynomial Density Hales Jewett theorem implies the Polynomial Szemeredi theorem. 

We consider the finitary statement of Sarkozy's theorem; that given any $0<\delta<1$, there would exist some integer $N(\delta)>0$ so that for all $n\geq N(\delta)$, any subset of $[n]$ with at least $\delta \cdot n$ many elements contains two numbers that differ by a non zero integer square. For the given $\delta$, consider the square of length $m_n=\lfloor \sqrt{2n^{2}+n} \rfloor$ and all the integers $n^{2}+i$ with $i\in A\subset [n]$ where $|A|\geq \delta\cdot n$. We can consider for each $i$, all possible sequences of 1's and 0's inside the $m_n \times m_n$ square grid so that the sum of the 1's is exactly $n^2+i$. It's apparent that for large enough $n$ we have a dense subset of $\{0,1\}^{m_n^2}$, and by \cref{conj:conj1} we would have two numbers of the form $n^2+i, n^{2}+i+|\gamma|^2$, where $\gamma$ is the size of the wildcard set, and where $i, i+|\gamma|^{2}$ both belong to $A$, which would establish Sarkozy's theorem.

In this paper, we show one can construct dense subsets of $\{0,1\}^{n^{2}}$ that have certain restrictions on the shape of the wildcard sets for the combinatorial line. It is easy to see, as should be clear from the discussion in the next section, that when we consider the set of all elements of $\{0,1\}^{n^{2}}$ with an even number of $1$'s, then upon symmetric differences, we also have a set with an even number of $1$s. This is a vector subspace of index 2 inside $\mathbb{F}_{2}^{n^{2}}$. The other coset is the set of elements with an odd number of $1$'s. One can thus consider either of these cosets which are dense subsets and the difference sets will always have an even nummber of elements. One can thus rule out wildcard sets of odd size in this case. Since the coloring Polynomial Hales Jewett theorem is true, the two cosets above can be considered to be a partition of $\{0,1\}^{n^2}$ into two colors, and at least one of them would have a combinatorial line and thus also a difference set of the form $\gamma \times \gamma$, and we would be forced to have $|\gamma|$ even. Because of the coset structure for subspaces in $\mathbb{F}_{2}^{n^{2}}$, once we find two elements whose difference set is of the square form $\gamma\times \gamma$, we also find the same in the other coset. This argument also extends in the obvious way for subspaces of any index $n\geq 3$.

Here we state the following theorem, the first case of which is already explained above, and the rest are proved in Section 3.  

\begin{theorem}\label{thm:mainthm1}

\begin{enumerate}
    \item[(i)] For large enough $n$, there exist subsets of $\{0,1\}^{n^2}$ containing exactly $2^{n^{2}-1}$ elements, thus of density $\delta=1/2$, so that the difference set of any two elements of this subset that constitute a combinatorial line, is of the form $\gamma \times \gamma$ with $|\gamma|$ is even.
    
    \item[(ii)] Given a positive integer $k$ and large enough $n$, and a set of elements $S=\{a_1,\dots,a_k\} \subset \{1,\dots,n\}$, there exist subspaces of $\{0,1\}^{n^2}$ of density $\delta=\frac{1}{2^{k}}$ so that the difference set of any two elements of this subspace that constitute a combinatorial line, does not contain any coordinate belonging to $S$.

    \item[(iii)] Given a positive integer $k$ and large enough even integer $n$, and a set of elements $S=\{a_1,\dots,a_k\} \subset \{1,\dots,n\}$, there exist subspaces of $\{0,1\}^{n^2}$ of density $\delta=\frac{1}{2^{\lfloor \frac{k}{2}\rfloor}}$ so that the difference set of any two elements of this subspace that constitute a combinatorial line, contains at least one pair of two consecutive elements of $S$.

    \item[(iv)] For any large enough $n\geq 4$, there exist subspaces of density $\delta=\frac{1}{4}$ so that the difference set corresponding to the combinatorial line contained in this subspace is of the form $\gamma \times \gamma$ with $\gamma\subset \{1,\dots,n\}$ so that $4 ||\gamma|$.
    
    \item[(v)] For any large enough $n\geq 4$, there exist subspaces of density $\delta=\frac{1}{2}$ so that the difference set corresponding to the combinatorial line contained in this subspace is of the form $\gamma \times \gamma$ with $\gamma\subset \{1,\dots,n\}$ so that $|\gamma|\equiv 0,-1 (\text{mod} \ 4)$.
    
    \item[(vi)] For any large enough $n\geq 4$, there exist subspaces of density $\delta=\frac{1}{2}$ so that the difference set corresponding to the combinatorial line contained in this subspace is of the form $\gamma \times \gamma$ with $\gamma\subset \{1,\dots,n\}$ and $\gamma$ is symmetric in the sense that whenever $j\in \gamma$, we have $(n+1-j) \in \gamma$, and further when $n$ is even, we have $4||\gamma|$.
    
    \item[(vii)] For each positive integer $t$, we have subspaces of density $\delta=\frac{1}{2}$, in $\{0,1\}^{n^{2}}$ for $n$ large enough, so that the difference set corresponding to the combinatorial line contained in this subspace is of the form $\gamma\times \gamma$ and if the initial and last elements of $\gamma$ are $a \ \text{and} \  b$ when arranged in increasing order, then $a+t, b-t\in \gamma$ as well.
    
\end{enumerate}
\end{theorem}    

\bigskip
The main theorem we prove in Section 4 is stated below:

\begin{theorem}\label{thm:mainthm2}
For any $\delta>0$, there exists $N(\delta)= \lceil \log_2(\frac{1}{\delta}+1)  \rceil$ such that for all $n\geq N(\delta)$, for any nonempty $\gamma_1\subset \{1,2,\dots,n\}$ we have for any subset $S$ of $\{0,1\}^{n^{2}}$ with $|S|\geq \delta\cdot 2^{n^{2}}$ , two elements $s_{1},s_{2}\in S$ such that their difference set is of the form $\gamma_{1} \times \gamma_{2}$ where $ \gamma_2\subset\{1,2,\dots,n\}$ is a non empty subset.   
\end{theorem}

In order to get a combinatorial line of the rectangular form, it will be enough to find two elements in $S$ such that their symmetric difference is of the rectangular form $\gamma_1 \times \gamma_2$, and the set of $1$'s in either of these two elements is a superset of this set $\gamma_1 \times \gamma_2$. This will likely require a large count of the number of pairs $(s_i,s_j)$ to force one pair to satisfy this property. This will be more involved than our proof where we obtain just one pair of elements with difference set of the rectangular form.

Following Furstenberg, Katznelson's original argument that showed that the multidimensional density Hales Jewett theorem follows from the density Hales Jewett theorem (Proposition 1.7 in \cite{Polymath}), we show that the natural conjectured multidimensional analog of the polynomial density Hales Jewett theorem follows from the conjectured quadratic density Hales Jewett theorem. 

\begin{theorem}\label{thm:mdqhj}($MQDHJ_{k}$) Let $\epsilon >0, k\in \mathbb{N}$. There exists $N=N(\epsilon,k)$ having the property that if $E\subset \{1,..,k\}^{N^{2}}$ with $|E|\geq \epsilon k^{N^{2}}$ then there exists a $d$ dimensional combinatorial subspace contained in $E$, of cardinality $k^{d}$, where each of the $d$ wildcard sets is of the form $\alpha_i \times \alpha_i$ where for $1\leq i\leq d$ , each of the $\alpha_i \subset \{ 1,2,..,N\}$ are distinct. 
\end{theorem}

\begin{proof}[Proof of Theorem \ref{thm:mdqhj}]
The induction is on the dimension $d$. Suppose we have proven $MQDHJ_{k}$ for $d-1$ and let $A\subset [k]^{N^{2}} $ have density $\epsilon>0$. Let $m=MQDHJ_{k}(\epsilon/2, d-1)$. For every string $z\in [k]^{N^{2}}$, consider the associated pair $z'=(z_{u},z_{l})$ where $z_{u}$ is the restriction of $z$ to the upper left $m^{2}$ square coordinates, and $z_{l}$ is the restriction to the lower $(N-m)^{2}$ square coordinates. Note that in going from $z$ to $z'$ we are neglecting the coordinates that do not belong to the union of the upper left $m^2$ square coordinates and the lower right $(N-m)^{2}$ many coordinates. 

Call a string $z_{l}$ in $[k]^{(N-m)^{2}}$ good, if the set $E_{z_{l}}= \{ z_{u} \in [k]^{m^2}: (z_{u},z_{l})\in E \}$ has density at least $\epsilon/2$ within $[k]^{m^{2}}$. It is not hard to see that the set of good strings has to have density at least $\epsilon/2$ within $[k]^{(N-m)^{2}}$. If not, we can bound from above the total number of elements in $E$ by $(\epsilon/2)\cdot k^{m^{2}}\cdot k^{(N-m)^{2}} + (\epsilon/2)\cdot k^{(N-m)^{2}}\cdot k^{m^{2}} =\epsilon \cdot k^{m^{2}+(N-m)^{2}}\leq \epsilon\cdot k^{N^{2}} $ which is a contradiction.

By induction, for any good $z_{l}$ the set $E_{z_{l}}$ contains a $(d-1)$ dimensional subspace. There are only a finite number of possible $(d-1)$ dimensional subspaces in $[k]^{m^{2}}$, and we call that number $T(m)$. By our choice of $m=MQDHJ_{k}(\epsilon/2,d-1)$, for any given good $z_{l}$, the set $E_{z_{l}}$ must have a $(d-1)$ dimensional subspace. 

For each $(d-1)$ dimensional subspace $\sigma$ in the upper left set of squares of cardinality $k^{m^{2}}$ construct the set $G_{\sigma}=\{ z_{l}:(x,z_{l})\in E  \ \  \forall x\in \sigma \}$ which is a subset of the lower right square of dimension $(N-m)\times (N-m)$. The union of the sets $G_{\sigma}$ over all possible $(d-1)$-dimensional subspaces covers the set of good $z_{l}$'s. Since the set of good strings has a density at least $\epsilon/2$ within $[k]^{(N-m)^{2}}$, at least one of the sets $G_{\sigma_{0}}$ has density at least $\frac{\epsilon}{2T(m)}$ within $[k]^{(N-m)^{2}}$. Assuming the quadratic density Hales Jewett statement to be true, we can choose $N-m> QDHJ_{k}(\frac{\epsilon}{2T(m)})$ and then we would get a 1-dimensional subspace $\lambda$ in the lower right set of squares, and then the subspace $\sigma_{0}\times \lambda$ would suffice for our purposes.
\end{proof}

\section{Symmetric differences and addition in the vector space $\mathbb{F}_{2}^{n^{2}}$}

Taking symmetric differences in our case is equivalent to the properties of addition in the vector space $\mathbb{F}_{2}^{n^{2}}$ over the field $\mathbb{F}_{2}$. Henceforth, we talk about addition in $\mathbb{F}_{2}^{n}$ in place of symmetric differences. We can get exactly $M=2^{m}-1$ many elements in a vector subspace upon addition in $\mathbb{F}_{2}^{n^{2}}$, when considering a vector subspace of $\mathbb{F}_{2}^{n^{2}}$ with $m$ basis vectors.

Let $m\leq n^{2}$. We consider any arbitrary $m$ elements $a_1,a_2,\dots, a_m$ among which there are no nontrivial relations through symmetric differences, this being possible since $\mathbb{F}_{2}^{n^2}$ is $n^{2}$ dimensional and $m\leq n^{2}$.

Considering all the sums from the basis set $\{a_1,\dots,a_m\}$, with coefficients in $\mathbb{F}_2$, we get exactly $M=(2^{m}-1)$ non zero elements. We enumerate these elements as $a_1,a_2,\dots,a_M$ keeping the first $m$ terms as before. This is a finite dimensional vector subspace with the sumset taking as many values as the number of elements in the subspace: for each $a_{k}, 1\leq k \leq M$, we have a unique chain of $2^{m-1}-1$ pairs of elements that add up to $a_k$; for $a_{k}$ and any other $a_l, l\neq k, 1\leq l\leq M$, the element $(a_{k}-a_{l})\in \mathbb{F}_{2}^{n^{2}}$ is added to $a_{l}$ to get $a_{k}$.\footnote{The question of finding dense subsets of $\mathbb{F}_{2}^{n}$ and checking whether the sumset has a large vector subspace, has been studied and there is an example of a set of density $1/4$ whose sumset does not contain a subspace of dimension $\sqrt{n}$. \cite{Green} } 

\section{Proof of \cref{thm:mainthm1}}
Within our $[n]^{2}$ grid, the first coordinate increases downwards, starting from 1, along a column when the second coordinate is fixed, and along a row the second coordinate increases, starting from 1, for a fixed value of the first coordinate. Given any $\delta=\frac{1}{2^{k}}$, in each case below we construct  specific subsets of $\{0,1\}^{n^2}$ of cardinality $\delta \cdot 2^{n^{2}}=2^{n^{2}-k}$. 

By a \textit{coordinate point} in $\{0,1\}^{n^2}$, we mean an element of the form $(x,y)\in \{0,1\}^{n^2}$ with $0\leq x,y\leq n$. Alternately, we also call such an element $(x,y)$ as a single element subset of $[n]^2$.

We recall that given two elements $a_1,a_2 \in \mathbb{F}^{n^2}_2$, the coordinate points in $\{0,1\}^{n^2}$ where they differ are precisely those coordinate points where the element  $(a_1 +a_2 )\subset \mathbb{F}^{n^2}_2$ takes the value $1$. 

We can construct $(n^{2}-k)$ many linearly independent subsets of $\mathbb{F}_{2}^{n^2}$ over $\mathbb{F}_{2}$, and get $2^{n^2 -k}-1$ many non-zero elements.

\begin{proof}[Proof of Theorem 6]

(i): This was already articulated in the introduction. We simply consider the set of elements of $\{0,1\}^{n^2}$ with an even number of $1$'s. Then, upon taking symmetric differences, we always have sets with an even number of $1$'s. Thus, any possible difference set also has an even number of elements.

(ii): In this case, consider a vector subspace consisting of all the coordinate points in $\{0,1\}^{n^2}$, except for the diagonal points $\{a_1,a_1\},\{a_2,a_2\},\dots,\{a_k,a_k\}$. In this case, it is clear that one cannot construct a combinatorial line which contains any one of the points from the set $S$.

(iii): For the case where $k=2k'$ is even, we pair off the elements of the diagonal as $\big\{(a_1,a_1),(a_2,a_2)\}\dots,\{(a_{2k'-1},a_{2k'-1}),(a_{2k'},a_{2k'})\}$. We consider a vector subspace where each coordinate point in  $\{0,1\}^{n^2}$ other than the ones considered already in the pairing above, are taken as basis elements. Thus, in the process of taking symmetric differences, we will not be able to construct a wildcard set $\gamma\times \gamma$ avoiding at least one pair of consecutive elements of $S$ . We have clearly constructed a subspace with $2^{n^2-k'}$ elements, and thus a subspace of density $\frac{1}{2^{\lfloor \frac{k}{2} \rfloor}}$. In the case where $k=2k'+1$ is odd, we again pair off the following elements of the diagonal set: $\big\{(a_1,a_1),(a_2,a_2)\}\dots,\{(a_{2k'-1},a_{2k'-1}),(a_{2k'},a_{2k'})\}$ and exclude the coordinate point $\{(a_{2k' +1},a_{2k' +1})\}$, and the result follows with this subspace also having density $\frac{1}{2^{\lfloor \frac{k}{2}\rfloor}}$.

(iv) Consider the set of all elements of $\{0,1\}^{n^2}$ that have an even number of elements in the diagonal; which is clearly a set of density $\frac{1}{2}$. Within this set, consider further the subset that has an even number of elements in the upper triangular part of the $[n]^{2}$ grid. Thus this set has a density of $\frac{1}{4}$ in $\{0,1\}^{n^2}$. In order to get a wildcard set, we would only get an even number of elements on the diagonal. Further we could only have an even number of elements in the upper triangular part of the grid. If we considered $4k+2$ many elements on the diagonal, then we would have to consider $1+2+\dots+(4k+1)=(2k+1)(4k+1)$ many points in the upper triangular grid, and this is a contradiction. Thus the only possibility is to get wildcards sets $\gamma \times \gamma$ with $|\gamma|$ divisible by 4. 

(v) Consider the basis set in this case, which has the following three types of elements: firstly the elements $\{(x,y)\}, \forall x,y\in [n]$ where $ x>y$, then the elements of the form $\{(t,t),(t+1,t+1)\}$ with $t\in [n-1]$, and further the elements of the form $\{(x,y),(x,y+1)\}$ when $x\leq y$ and when $(y+1)\leq n$. Consider any wildcard set of form $\gamma\times \gamma$ with $|\gamma|=k$ that is obtained, and enumerate the elements in $\gamma$ in increasing order as $\{t_1,\dots,t_k\}$. In order to include the point $(t_k,t_k)$, we have to include all the 2-element basis sets on the diagonal; $\{(t_{k}-1,t_{k}-1),(t_k,t_k)\}+\{(t_{k}-2,t_{k}-2),(t_{k}-1,t_{k}-1)\}+\dots+\{(t_{k-1},t_{k-1}),(t_{k-1}+1,t_{k-1}+1)\}=\{(t_{k-1},t_{k-1}),(t_k,t_k)\}$. Further, to include the point $(t_{k-1},t_{k})$ the only option we have is to include the basis sets $\{(t_{k-1},t_{k-1}),(t_{k-1},t_{k-1}+1)\},\dots, \{(t_{k-1},t_{k}-1),(t_{k-1},t_{k})\}$ which add up to the set $\{(t_{k-1},t_{k-1}),(t_{k-1},t_{k})\}$.

 Till now the point $(t_{k-1},t_{k-1})$ was added exactly twice and thus has been excluded. The only option we now have is to further include all the two element diagonal elements starting from $(t_{k-2},t_{k-2})$ till $(t_{k-1},t_{k-1})$, (in a  way similar to how we included all the elements from $(t_{k-1},t_{k-1})$ till $(t_{k},t_{k})$ in the previous paragraph). We would now have to include the elements $\{(t_{k-2},t_{k-1}),(t_{k-2},t_k)\}$ by adding up a suitable number of basis elements in the upper triangular region. This would be permissible as a wildcard set of shape $\gamma\times \gamma$ with $|\gamma|=3$. We can extend this to possibly have $|\gamma|=4$ where we have four further points to account for in the same row with the first coordinate $t_{k-3}$, which can be done obviously as before, and this is the only way to account for the point $(t_{k-3},t_{k-3})$ (If we accounted for the point $(t_{k-3},t_{k-3})$ by covering it with the set of the form $\{(t_{k-3}-1),(t_{k-3}-1),(t_{k-3},t_{k-3})\}$ and so forth along the diagonal, then we would be left to cover three elements in the upper triangular part in the row $t_{k-3}$ which is impossible with our basis elements). So in this way, we can get a wildcard set of size either $3$ or $4$. It is now clear that we can extend the argument to get wildcard sets only of size $4k$ or $4k-1$ for any $k\geq 2$. We have one element basis sets in the lower triangular part of the grid, and so we can add each required coordinate point. It is also obvious that we have exactly $2^{n^{2}-1}$ many elements in this subspace, and thus is a subspace of density $\frac{1}{2}$.
 
 (vi) In this case, we choose for the basis, all the permissible two element sets of the form $\{(a,b),(a-1,b+1)\}$ with $a+b=n+1$, which are all the two element subsets of the ``reverse" diagonal. Every other coordinate point is considered as a basis element. This set has density $\frac{1}{2}$. It is clear that we must attain at least one element from this 'reverse diagonal' set when taking linear combinations of these basis sets, and thus it is clear that whenever $j\in \gamma$ we must have that $(n+1-j)\in \gamma$ as well. Further when $n$ is even, and thus $n+1$ is odd, we have exact pairings of distinct elements and there are only an even number of elements of this reverse diagonal one can cover, and thus we must have $4|n$.
 
 (vii) In this case, we consider the consecutive two element basis sets of the type we considered in the previous example, but on the diagonal line in the upper triangular part of the grid parametrized by the line $y=x+t$, and every other single point on the grid is also taken as a basis point. It is clear that on this diagonal line parametrized by $y=x+t$ we must have attained the coordinate point $(a,a+t)$ where $a$ is the least point of $\gamma$ , as well as for the largest element $b\in \gamma$, the point $(b-t,b)$ being covered. This subspace also clearly has $2^{n^{2}-1}$ elements and thus is a subset of density $\frac{1}{2}$.

\end{proof}

\section{Rectangular and square wildcard sets}

\begin{proof}[Proof of \cref{thm:mainthm2}]
Observe that $(\gamma_a \times \gamma_b)+(\gamma_a \times \gamma_c)=\gamma_a \times (\gamma_b +\gamma_c)$ where addition is in $\mathbb{F}_{2}^{n^2}$.

Consider for any fixed $\gamma$, all the nonempty sets of the form $\gamma_{i}:=\gamma\times t_i$ with $t_i\subset \gamma$. There are $2^{n}-1$ many such sets. Suppose we have a subset $S\subset \{0,1\}^{n^2}$ with density $\delta$, whose elements we enumerate as $s_1.s_2,\dots,s_{|S|}$. If there exist some $\gamma_i, \gamma_j$ such that $\gamma_i+s_{t}=b, \gamma_j +s_{l}=b$ for the same fixed $b\in \{0,1\}^{n^2}$, then $\gamma_i+\gamma_j=\gamma\times (t_i+t_j)=s_{l}+s_{t}$ (Clearly $\gamma\times (t_i+t_j)$ is of the form $\gamma\times (t_k)$ for some nonzero $t_k\subset \gamma$). Such a thing is forced for values of $n \geq \lceil \log_{2}\big( \frac{1}{\delta} +1  \big)\rceil$: consider the set of all pairs of elements $(\gamma_i,s_{t})$ of which there are $(2^{n}-1)\cdot \delta\cdot 2^{n^2}$ many. If for any pair $(\gamma_{i},s_{t})$ we have $\gamma_i +s_t =s_j$ for some $j\in [S]$, we are done since then $s_i +s_j= \gamma_i$. Otherwise each of these pairs sum to some element in $\{0,1\}^{n^{2}}\setminus S $, and there are $(1-\delta)\cdot 2^{n^2}$ many such elements. Thus there has to be some fixed $b$ to which two pairs sum to, in which case again we are done as before.
\end{proof}

We could specialise the above argument where the rectangular difference set obtained actually has a form $\gamma_1 \times \gamma_2$ with $|\gamma_1|=|\gamma_2|$ but where the sets $\gamma_1$ and $\gamma_2$ are not necessarily equal. That is the content of the first part of the following theorem\footnote{After the completion of this work, it came to the notice of the author that this theorem has also been mentioned in a different language, in the end of the blog post of \cite{Gowersblog}.}. Also, if instead of square wildcard sets, we require wildcard sets that are the disjoint unions of two square wildcard sets, i.e. of the form $\gamma_1\times\gamma_1 +\gamma_2\times\gamma_2$ with $\gamma_1,\gamma_2\subset [n]$, then again the proof follows with essentially the same basic argument as above.

\begin{theorem}\label{thm:mainthm3}
(i)For any $\delta>0$ and even integer $l=2l_1>0$, there exists $N(\delta)= \lceil \frac{l}{2\delta}  \rceil$ such that for all $n\geq N(\delta)$, for any nonempty $\gamma_1\subset [n]$ with $|\gamma_1|=l$, we have for any subset $S$ of $\{0,1\}^{n^{2}}$ with $|S|\geq \delta\cdot 2^{n^{2}}$, two elements $s_{1},s_{2}\in S$ such that their difference set is of the form $\gamma_{1} \times \gamma_{2}$ with $|\gamma_2|=l$ as well.

(ii) For any $\delta>0$ and integer $l>0$, there exists $N(\delta)= \lceil \frac{l}{\delta}  \rceil$ such that for all $n\geq N(\delta)$, we have for any subset $S$ of $\{0,1\}^{n^{2}}$ with $|S|\geq \delta\cdot 2^{n^{2}}$, two elements $s_{1},s_{2}\in S$ such that their difference set is of the form $\gamma_1\times\gamma_1 +\gamma_2\times\gamma_2$ with $\gamma_1,\gamma_2\subset [n]$ with $|\gamma_1|=|\gamma_2|=l$.
\end{theorem}

\begin{proof}[Proof of \cref{thm:mainthm3}]

(i) Consider all pairs of elements of the form $(\gamma_1 \times \Tilde{\gamma},s)$ where $\Tilde{\gamma}$ are all the $l_1$-element subsets of $[n]$ of the form $\{tl_1+1,\dots,(t+1)l_1\}$ for $0\leq t<\frac{n}{l_1}$, and $s\in S$, where $n\geq N(\delta)$ and $S$ is as in the statement of the theorem. Clearly we have $(\frac{n}{l_1})\cdot\delta\cdot 2^{n^{2}}$ many distinct such pairs. In this case, when $\frac{n\delta}{l_1}\geq 1$, i.e. when $n\geq N(\delta)$, we must have that two distinct pair sums are the same, and then it is clear that we will find $s_1+s_2=\gamma_1\times \gamma_2$ with $\gamma_2=\Tilde{\gamma_1}+\Tilde{\gamma_2}$ with both $\Tilde{\gamma_1},\Tilde{\gamma_2}$ consecutive $l_1$ element subsets.

(ii) In this case, we consider pairs $(\gamma\times \gamma,s)$ with $\gamma\subset [n]$ of the form $\{tl_1+1,\dots,(t+1)l_1\}$ for $0\leq t<\frac{n}{l}$, and $s\in S$ and the argument proceeds as above.
\end{proof}

The stronger version of the first part of the above theorem would say that dense subsets of $\{0,1\}^{n^{2}}$ would contain a combinatorial line such that their corresponding difference set is of the form $\gamma_1\times \gamma_2$ with $|\gamma_1|=|\gamma_2|=k$ for any a-priori given positive integer $k$. Following the argument after Conjecture 2 in the introduction, this would also imply Sarkozy's theorem, but notably this will not generalize the coloring version of the quadratic Hales Jewett theorem where we must require $\gamma_1=\gamma_2$ as subsets of $[n]$

\section{Further questions.}

We state three natural conjectures below of which the first two are weaker versions of the main density quadratic Hales Jewett conjecture.

\textit{Conjecture 1:} For any $0<\delta<1$, there exists $N(\delta)$ such that for any $n\geq N(\delta)$, for any subset $S\subset \{0,1\}^{n^{2}}$ we must have two elements $s_1,s_2\in S$ such that their difference set is of the form $\gamma\times \gamma$ with $\gamma \subset [n]$.  

\bigskip

\textit{Conjecture 2:} For any $0<\delta<1$, there exists $N(\delta)$ such that for any $n\geq N(\delta)$, for any subset $S\subset \{0,1\}^{n^{2}}$ we must have two elements $s_1,s_2\in S$ such that their difference set is of the form $\gamma_1\times \gamma_2$ with $\gamma_1,\gamma_2 \subset [n]$ and further $\gamma_1\times \gamma_2$ is the wildcard set of a corresponding combinatorial line .  

\bigskip

We define a restriction on $[n]$, as the property that in any nonempty subset of $[n]$, a given pattern or sequence of elements is forbidden from appearing. As an example, say that a subset $S\subset [n]$ satisfies a specific restriction of there being no single element in isolation in the subset $S\subset [n]$, which means there cannot exist any $a\in S$ such that $a-1\notin S,a+1\notin S$. Another example is where we have a restriction that the cardinality of the given subset $S\subset [n]$ is not divisible by any given integer.
Each part in Theorem 6 involves a particular restriction.

\textit{Conjecture 3:} Given any specific restriction, there exist dense vector subspaces of $\{0,1\}^{n^{2}}$ for any large enough $n$, so that the wildcard set $\gamma\subset [n]$ corresponding to a combinatorial line satisfies the restriction.

\bigskip

In fact, one would expect to extend Conjecture 1 and for any given $\delta$, quantify a large enough proportion of pairs of elements so that their difference set is of the square form. Out of that quantifiably large proportion of pairs one would expect to find a pair which actually forms a combinatorial line. 
 
Conjecture 2 is a strengthening of Theorem 7 where we require to find a pair of elements forming a combinatorial line with the rectangular wildcard set.

In case of just the square wildcard sets of the form $\gamma \times \gamma$, we do not have the simple case of two square wildcard sets adding up to a square wildcard set, as was the case in the corresponding problem of the rectangular wildcard sets in Theorems 7 and 9. In this case, for any set $\gamma \subset \{1,2,\dots n\}$ with $|\gamma|\geq 3$, we would need to add up several more sets to get zero. For this set $\gamma$, consider the set of all nonempty subsets $\mathscr{P}_{\gamma}$ of $\gamma$. Consider the sets $\{a\times a : a \subset \mathscr{P}_{\gamma} \}$. Upon adding all the elements of this set we would get 0: every $(i,i)$ inside $\gamma \times \gamma$ gets counted exactly $(2^{|\gamma |-1}-1)+1$ many times, and every pair $(i,j)\in \gamma \times \gamma$ with $i\neq j$ gets counted $(2^{|\gamma|-2}-1)+1$ many times, both even for $|\gamma|>3$. One can also consider for some disjoint $\gamma_1,\gamma_2 \subset \{1,2,\dots n\}$, the sets $\{(a\cup \gamma_{2})\times (a\cup \gamma_2): a\subset \mathcal{P}_{\gamma_1}\}$, where $\mathcal{P}_{\gamma_1}$ is the set of all subsets (including the empty set) of $\gamma_{1}$, with again $|\gamma_{1}|\geq 3$. In this case too, upon adding all the elements we would get $0$. Building on this would be a significantly more challenging problem which remains for future exploration.

Conjecture 3 would likely involve extending the methods of Theorem 6 to find subspaces that force any possible given restriction on the wildcard sets. It would be interesting to classify the kinds of restrictions that are actually not permissible.

\section{Acknowledgements}

The author is thankful to Daniel Glasscock for pointing out the quadratic density Hales Jewett conjecture to the author, and also to Tim Austin for useful feedback on an earlier version of the manuscript. The author also acknowledges useful discussions with John Griesmer, and his pointing out an error in an earlier draft.

\end{document}